Fermi National Laboratory


Accelerator Update for September 7, 2005 - September 9, 2005

Wednesday, September 7
The day shift began with Operations monitoring store 4371 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, and with MiniBooNE and NuMI taking beam.

The Meson MCenter experiment resumed taking beam at 8:20 AM.

A RF expert began a Main Injector (MI) RF study at 9:33 AM.

TeV experts began an end-of-store study at 2:42 PM.

Pbar experts began a study period around 2 PM.

At 3:15 PM, the Meson MS2 LCW system alarmed. An expert investigated and found that a cooling tower had tripped off.

Operations terminated store 4371 at 3:34 PM. Operations prepared CDF and D0 for collision hall accesses and the TeV tunnel for access.

At 3:39 PM, Operations reported that the Recycler lost 50mA of antiprotons. Later, an expert reported that a bug in some recently installed code had caused the emittances to blow up.

At 4:59 PM, a TeV expert reported a burning smell coming from B0Q5 power supply. Power supply experts investigated and found a bad contactor. They began repairs.

Pbar studies ended at 8:08 PM and Pbar resumed stacking twenty minutes later.

Operations put the Recycler into shot setup at 8:40 PM to stash antiprotons.

D0 came out of access at 9:27 PM.

The Recycler stash ended and Pbar resumed stacking at 9:39 PM. Operations reported that the efficiency for the last three transfers was very poor due to a low level RF problem. Experts will investigate further on Thursday morning.

Power supply experts reported at 9:26 PM that they had completed their repairs on B0Q5.

Thursday, September 8
The midnight shift began with Operations monitoring the Recycler stash, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

Recycler experts reported at 5:26 AM that the low level RF system had to be rebooted and, a bit later, was successfully accomplished.

Recycler began preparing to stash antiprotons at 6:03 AM. The stash was completed and stacking resumed at 6:42 AM.

At 8:25 AM, a power supply expert reported that he had to do more work on B0Q5.

A RF expert began an upgrade of the MIRF stations at 10:27 AM. This upgrade will get done one station at a time as time permits.

A MI expert began a coalescing study at 10:28 AM.

Operations put the Tevatron (TeV) into shot setup at 12:14 PM.

Operations established combination store 4374 at 2:24 PM with an initial luminosity of 110.39E30. Pbar resumed stacking. MCenter had gone into access at noon.

The Recycler began preparing to stash antiprotons 6:40 PM. Operators completed the stash and Pbar resumed stacking at 7:39 PM.

Operations established beam to MCenter at 7:46 PM.

At 8:10 PM, MIRF stations 2, 5, & 13 tripped off. They reset okay. The NuMI beam permit tripped off due to the MIRF trip. It reset okay.

Operations reported at 8:34 PM that the Recycler Electron Cooling (ECool) was off for the night.

At 9:07 PM, Operations reported that a TeV sector B11 separator sparked and caused a large drop in luminosity.

Operations terminated store 4374 at 9:44 PM.

At 10:21 PM, MIRF2, 3, & 13 tripped off. They reset okay.

TeV experts began a study period around 10:10 PM.

Friday, September 9
The midnight shift began with Operations monitoring the stash, with TeV experts conducting studies, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 12:18 AM, MIRF13 & 14 tripped off. These stations plus several others tripped off for the first of many times over the next three and a half hours. Each station reset okay.

TeV experts reported that they had completed a collimator study at 2:29 AM and were beginning an orbit study.

At 4 AM, MIRF13 tripped off. It took ten minutes before operators could get the station to reset.

TeV studies ended at 4:22 AM.

Operations put the TeV through a dry squeeze at 4:42 AM.

Operation began preparing the TeV for shot setup around 5:07 AM.

At 7:52 AM, Operations reported that various problems with the TeV tune up, proton qualifiers, and coalescing were delaying shot setup.

The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 9/9/2005   email Fermilab